a= 0.05 test the claim that u>16 using P value n= 50 x=16.3 o= 1.2 find p value.
Found 2 solutions by ewatrrr, stanbon:
Answer by ewatrrr(24785)

Ho: u = 16
Ha: u >16 (claim)
.......
sample of 50: xbar =16.3 o= 1.2minutes.
t(16.3) = .3/(1.2/sqrt(50)) = 1.7678
p-value = tcdf (1.7678, 100, 49) = .0417
......
If Using z-value: normalcdf(1.7678, 100) = .0385 (Same Results)
.......
confidence level .05
.04 < .05 0r .0385 < .05
Reject Ho. Test results support the claim u >16
......
re TY
mean of 1400 and a standard deviation of 200.
A college will only take applicants in the top 2.5% or higher.
What is the lowest score they would need on the entrance exam
200(invNorm(.975) + 1400 = 200(1.96) + 1400 = 1792, lowest score

Answer by stanbon(75887)

a= 0.05 test the claim that u>16 using P value n= 50 x=16.3 o= 1.2 find p value.
------
z(16.3) = (16.3-16)/1.2 = 0.3/1.2 = 0.25
----
p-value = P(z > 0.25) = normalcdf(0.25,100) = 0.40
-----
Note: You may have meant that x-bar = 16.3 in which case
z(16.3) = (16.3-16)/(1.2/sqrt(50)) = 1.77
Then p-value = P(z > 1.77) = normalcdf(1.77,100) = 0.038
------------
Cheers,
Stan H.
------------